### v1

@@ 6 Summary and Conclusions
In this study, we combine a simulated Milky Way (MW) analogue for HESTIA suite of high-resolution galaxy formation simulations (Libeskind et al., 2020), and the New Generation Planet Population Synthesis dataset from the Bern model for planet formation and evolution (Emsenhuber et al., 2021a), in order to study the planet demographics throughout the galaxy. Using the stellar ages and metallicities provided by HESTIA, we are able to map the occurrence rate of various planet types as a function of galactic environment, and quantify variations in the frequency of Earth-like planets, Super-Earths, Neptunians, and Giant planets (defined through their planetary masses). Our key findings are as follows:
@@
The study of planet formation as a function of galactic environment is still in its infancy. Currently, transit and RV surveys are restricted mostly to stars within 1 kpc of the Sun. Future exoplanet surveys will however increase our understanding of planet formation and the effects of the environment. The upcoming PLATO mission (Rauer et al., 2014) will observe around 200’000 stars covering a large area of the sky, and is likely to detect a sample of exoplanets around thick disc stars. In addition, the Nancy Grace Roman Space Telescope (Spergel et al., 2015) is expected to find a statistical sample of around 1400 planets in the galactic bulge (Penny et al., 2019) through microlensing. Understanding the exoplanet demographics in different environments will yield new insights into processes involved in planet formation and evolution.
@@
Code and Data Availability
The code used for the analysis in this work has been compiled into a python module skaro and can be found at github.com/ChrisBoettner/skaro. The NGPPS data is provided by Emsenhuber et al. (2021a) and is available at dace.unige.ch. Access to the HESTIA data (Libeskind et al., 2020) was kindly provided by the CLUES collaboration (clues-project.org), and can be requested from them. The tables and further data presented in this work are available upon request to boettner@astro.rug.nl.
@@ Acknowledgements.
CB thanks the Young Academy Groningen for their generous support through an interdisciplinary PhD fellowship. PD and MT acknowledge support from the NWO grant 016.VIDI.189.162 (“ODIN”). PD warmly thanks the European Commission’s and University of Groningen’s CO-FUND Rosalind Franklin program. K.R. is grateful for support from the UK STFC via grant ST/V000594/1. We want to thank Alexandre Emsenhuber for providing us with the Monte Carlo variable dataset used for the original NGPPS sample, and the Data & Analysis Center for Exoplanets (dace.unige.ch) for a seamless access to the NGPPS and observational exoplanet data.References
@@
Bulge |
Thin Disc |
Thick Disc |
Halo |
1 |
Earth |
3.6 |
3.9 |
5.5 |
5.4 |
Super-Earth |
3.9 |
4.1 |
2.7 |
2.4 |
Neptunian |
0.68 |
0.60 |
0.13 |
0.11 |
Sub-Giant |
0.146 |
0.135 |
0.019 |
0.014 |
Giant |
0.407 |
0.313 |
0.021 |
0.018 |
0.5 |
Earth |
4.6 |
5.0 |
4.8 |
4.3 |
Super-Earth |
3.2 |
3.5 |
1.0 |
0.8 |
Neptunian |
0.34 |
0.38 |
0.04 |
0.03 |
Sub-Giant |
0.084 |
0.075 |
0.002 |
0.003 |
Giant |
0.022 |
0.020 |
0.002 |
0.001 |
0.3 |
Earth |
5.0 |
5.1 |
4.5 |
4.4 |
Super-Earth |
2.1 |
2.1 |
1.7 |
1.7 |
Neptunian |
0.11 |
0.12 |
0.08 |
0.07 |
Sub-Giant |
0.019 |
0.016 |
0.004 |
0.002 |
Giant |
0.000 |
0.000 |
0.000 |
0.000 |

### v2

@@ 6 Summary and conclusions
In this study, we combined a simulated MW analogue for HESTIA suite of high-resolution galaxy formation simulations (Libeskind et al., 2020) and the NGPPS dataset from the Bern model for planet formation and evolution (Emsenhuber et al., 2021a) to study the planet demographics throughout the galaxy. Using the stellar ages and metallicities provided by HESTIA, we were able to map the occurrence rate of various planet types as a function of galactic environment and quantify variations in the frequency of Earth-like planets, Super-Earths, Neptunians, and giant planets (defined through their planetary masses). Our key findings are as follows:
@@
The study of planet formation as a function of galactic environment is still in its infancy. Currently, transit and RV surveys are restricted mostly to stars within 1 kpc of the Sun. Future exoplanet surveys will however increase our understanding of planet formation and the effects of the environment. The upcoming PLATO mission (Rauer et al., 2014) will observe around 200 000 stars covering a large area of the sky and is likely to detect a sample of exoplanets around thick-disc stars. In addition, the Nancy Grace Roman Space Telescope (Spergel et al., 2015) is expected to find a statistical sample of around 1400 planets in the galactic bulge (Penny et al., 2019) via microlensing. Understanding the exoplanet demographics in different environments will yield new insights into processes involved in planet formation and evolution, while helping improve planet formation models.
@@ Acknowledgements.
CB thanks the Young Academy Groningen for their generous support through an interdisciplinary PhD fellowship. PD and MT acknowledge support from the NWO grant 016.VIDI.189.162 (“ODIN”). PD warmly thanks the European Commission’s and University of Groningen’s CO-FUND Rosalind Franklin program. K.R. is grateful for support from the UK STFC via grant ST/V000594/1. We want to thank Alexandre Emsenhuber for providing us with the Monte Carlo variable dataset used for the original NGPPS sample, and the Data & Analysis Center for Exoplanets (dace.unige.ch) for a seamless access to the NGPPS and observational exoplanet data. Access to the HESTIA data (Libeskind et al., 2020) was kindly provided by the CLUES collaboration (clues-project.org), and can be requested from them. The tables and further data presented in this work are available upon request to boettner@astro.rug.nl. We also would like to thank the referee for insightful feedback, which helped to improve this work.References
@@
- Schib et al. (2021) Schib, O., Mordasini, C., Wenger, N., Marleau, G. D., & Helled, R. 2021, Astronomy and Astrophysics, 645, A43
- Schlaufman (2018) Schlaufman, K. C. 2018, The Astrophysical Journal, 853, 37
@@
| Bulge | Thin Disc | Thick Disc | Halo | ||
| 1 | Earth | 3.6 | 3.9 | 5.5 | 5.4 |
| Super-Earth | 3.9 | 4.1 | 2.7 | 2.4 | |
| Neptunian | 0.68 | 0.60 | 0.13 | 0.11 | |
| Sub-giant | 0.146 | 0.135 | 0.019 | 0.014 | |
| Giant | 0.407 | 0.313 | 0.021 | 0.018 | |
| 0.5 | Earth | 4.6 | 5.0 | 4.8 | 4.3 |
| Super-Earth | 3.2 | 3.5 | 1.0 | 0.8 | |
| Neptunian | 0.34 | 0.38 | 0.04 | 0.03 | |
| Sub-giant | 0.084 | 0.075 | 0.002 | 0.003 | |
| Giant | 0.022 | 0.020 | 0.002 | 0.001 | |
| 0.3 | Earth | 5.0 | 5.1 | 4.5 | 4.4 |
| Super-Earth | 2.1 | 2.1 | 1.7 | 1.7 | |
| Neptunian | 0.11 | 0.12 | 0.08 | 0.07 | |
| Sub-giant | 0.019 | 0.016 | 0.004 | 0.002 | |
| Giant | 0.000 | 0.000 | 0.000 | 0.000 |